OpenStudy (anonymous):

Find two consecutive even integers whose product is equal to 10,200

jimthompson5910 (jim_thompson5910):

one even number is 2x for some integer x the next consecutive even integer is 2x+2

jimthompson5910 (jim_thompson5910):

Their product is 10200, so... (2x)*(2x+2) = 10200

OpenStudy (anonymous):

meaning................

jimthompson5910 (jim_thompson5910):

Distribute to get (2x)*(2x+2) = 10200 2x*2x+2x*2 = 10200 4x^2 + 4x = 10200

OpenStudy (stephstar101):

x(x+2)=10200 <- X is one even number, x+2 is the next consecutive even number x^2+x=10200 <- multiplied x and x+2 from previous step x^2+x-10200=0 <- subtract 10200 from both sides from previous step (x-102)(x-100)=0 <-factor them out The last step is the answer! Find both Xs!

jimthompson5910 (jim_thompson5910):

Now solve for x

OpenStudy (anonymous):

got it thanks guys!